1.
A(n) ________________ contains a whole number part and a fraction part
[1] Proper fraction
[2] Complex fraction
[3] Improper fraction
[4] Mixed number

2.The following are ordered from to greatest is.
1.1/10;4/9;3/5
2.4/9;1/10;3/5
3.3/5;1/10;4/9
4.4/9;3/5;1/10

3.John covers 2/3
of a journey by car 1/4
of the journey by bicycle, and walks the rest of the way. What part of does he walk
1.3/7 2.1/8 3.1/12 4.3/10
1
Expert's answer
2020-07-22T18:14:36-0400
  1. [4] Mixed number
  2. [1]1/10;4/9;3/5
  3. [3]1/12


Details

2.

We've three different numbers with different denominators - we can't compare them. To compare them, we need to have the same denominator.

Now, "LCM(10, 9, 5) = 90". We need to have this denominator. Thus,


"\frac{1}{10} = \frac{9}{90}, \frac{4}{9} = \frac{40}{90}, \frac{3}{5} = \frac{60}{90}"

Thus we can compare them with respect to their numerators.


3.

We will take the complete work as 1. John covers "\frac23" distance by car and "\frac14" by bicycle.

So total distance covered this way "= \frac23 + \frac14 = \frac{8}{12} + \frac{3}{12}=\frac{8+3}{12}=\frac{11}{12}".

Distance left "= 1 - \frac{11}{12} = \frac{12}{12} - \frac{11}{12} = \frac{12-11}{12}=\frac{1}{12}".